ELECTRONIC SPECTROSCOPY OF WC FROM 17,500 to 24,000 cm^{−1}

\underline{A. W. SMITH}, S. M. SICKAFOOSE, and M. D. MORSE, Department of Chemistry, University of Utah, Salt Lake City, UT, 84112.

Resonant two-photon ionization spectroscopy has been used to study the diatomic molecule WC. A low resolution scan revealed a five member vibrational progression beginning at 17 585 cm^{−1} with the ν_{00} transition. Analysis of this progression yielded a vibrational frequency of ω′_{e}(^{184}W^{12}C) = 752.6(4.9) cm^{−1} and a bond length of r′_{e}(^{184}W^{12}C) = 1.747(4) Å. In addition, several unassigned bands were rotationally resolved. Interestingly, all of the observed excited states have Ω′ = 2. All of the rotationally resolved transitions were fit simultaneously to produce the best possible fit of the ground state. Assignment of these bands confirmed a ground state of ^{3}Δ_{1} from a 14σ^{2}8π^{4}15σ^{2}4δ^{1}16σ^{1} configuration and determined the ground state as r′′_{0}(^{184}W^{12}C) = 1.7143(2) Å. Dispersed fluorescence studies to elucidate the ground and low-lying excited states will also be reported. These results on WC are compared to the results of studies on MoC and other transition metal carbides.